A battery of e.m.f E and negligible internal resistance R1 and R2 connected in parallel. Calculate the current through R2 in terms of R1

1
Expert's answer
2021-11-04T18:04:21-0400

"R_1,R_2" Connected parallel combination

"R_{eq}=\frac{{R_1}{R_2}}{R_1+R_2}"

Current flow "R_1" Resistance


"I_1=\frac{E}{R_1}"

Current flow in"R_2" Resistance

"I_2=\frac{E}{R_2}"